# HERACLES++: a multi-dimensional Eulerian code for exascale computing

### v1

@@
HERACLES++: a multi-dimensional Eulerian code for exascale computing
@@
Numerical simulations of multidimensional astrophysical fluids present considerable challenges. However, the development of exascale computing has significantly enhanced computational capabilities, motivating the development of new codes that can take full advantage of these resources. In this article, we introduce HERACLES++, a new hydrodynamics code with high portability, optimized for exascale machines with different architectures and running efficiently both on CPUs and GPUs. The code is Eulerian and employs a Godunov finite-volume method to solve the hydrodynamics equations, which ensures accuracy in capturing shocks and discontinuities. It includes different Riemann solvers, equations of state, and gravity solvers. It works in Cartesian and spherical coordinates, either in 1-D, 2-D, or 3-D, and uses passive scalars to handle gases with several species. The code accepts a user-supplied heating or cooling term to treat a variety of astrophysical contexts. In addition to the usual series of benchmarking tests, we use HERACLES++ to simulate the propagation of a supernova shock in a red-supergiant star envelope, from minutes after core collapse until shock emergence. In 1-D, the results from HERACLES++ are in agreement with those of V1D for the same configuration. In 3-D, the Rayleigh-Taylor instability develops and modifies the 1-D picture by introducing density and composition fluctuations as well as turbulence. The focus on a wedge, rather than the full solid angle, and the ability of running HERACLES++ with a large number of GPUs allow for long-term simulations of 3-D supernova ejecta with a sub-degree resolution. Future development is to extend HERACLES++ to a radiation-hydrodynamics code.
@@ 1 Introduction
Massive stars end their lives in the gravitational collapse of their degenerate core, the formation of a compact object, and possibly followed by an energetic explosion. This phenomenon is dubbed a core-collapse supernova (SN; Baade & Zwicky 1934). Most of the binding energy released by the collapsing envelope is emitted as neutrinos (Arnett et al., 1989). If revived, the SN shock born at core bounce propagates through the progenitor envelope and its highly stratified composition structure. During its expansion through the progenitor envelope, hydrodynamical instabilities grow, leading to chemical mixing on small and large scales. Spatially resolved SN remnants provide direct evidence for this chemical mixing, as well as the heterogeneous density structure of SN ejecta, with well documented examples such as SN 1987A or Cas A (Fesen et al., 2006; Abellán et al., 2017; Milisavljevic et al., 2024). In Type II SNe, which arise from the explosion of generally red-supergiant (RSG) stars, the main instability is Rayleigh-Taylor (Falk & Arnett, 1973). Its development is influenced by the post-shock neutrino-driven convection that develops as part of the explosion (Kifonidis et al., 2000, 2003), as well as the structured convective envelopes of these extended progenitor stars (Goldberg et al., 2022a, b). The nonlinear growth of the Rayleigh-Taylor instability in SN ejecta has been extensively studied (e.g., see Chevalier 1976) or more recently Duffell 2016. Most of these studies were initially focused on SN 1987A (Ebisuzaki et al., 1989; Mueller et al., 1991; Fryxell et al., 1991; Hachisu et al., 1994) and have been extended in recent years to Type II SNe in general (e.g., Wongwathanarat et al. 2015).
@@
When studying the development of instabilities in SN ejecta, the structure and chemical composition of the progenitor is critical. The Rayleigh-Taylor instability tends to develop at the H-He and the O-He interfaces (Herant & Benz, 1991; Mueller et al., 1991; Herant & Woosley, 1993), and the resulting structures take the form of fingers and mushrooms, altering the density and chemical profiles obtained in spherical symmetry (e.g., Wongwathanarat et al. 2015). The modeling of such fluids in 3-D is computationally intensive and it is therefore desirable to develop efficient and robust numerical tools to tackle these scientific problems.
@@
More and more computational capacities are available with the creation of new supercomputers and the development of new hardware. Exascale supercomputers operate with a massively parallel architecture and have a computing power exceeding $10^{18}$ floating-point operations per second (FLOPS). The exascale era offers the ability to solve large and complex problems, opening up new possibilities for research by combining unprecedented computing power and data analysis. Indeed, the computational resources rely less and less on CPUs and benefit more and more from GPUs. Practically, among the world’s fastest supercomputers, more than 75% of the available FLOPS are delivered by GPUs [^1] Computed from www.top500.org. GPUs have become a crucial driver of computational performance in high-performance computing (HPC) and different vendors (e.g., AMD, NVIDIA, Intel) are developing their own type of hardware, which can be used with their own programming language (e.g., HIP/ROCm, CUDA, OneApi). To fully exploit these possibilities, scientists must develop software and numerical approaches that use the benefits of exascale hardware and ensure portability across different architectures.
@@
To meet these challenges, and especially performance portability, we have developed the hydrodynamics code HERACLES++ using the Kokkos library (Trott et al., 2022). It allows for a unified C++ code portable to different architectures, offering a level of abstraction that enhances user-friendliness. The code uses Message Passing Interface (MPI) to deal with distributed memory parallelism and Kokkos for shared memory computations. With this approach, the code can efficiently use most of the CPUs and GPUs available on the market, including NVIDIAs and AMDs.
@@
The usage of Kokkos is slowly spreading and different codes have been recently developed using the Kokkos-C++ framework. For example, IDEFIX (Lesur et al., 2023) and ARK (Padioleau et al., 2019; Bourgeois et al., 2024) are fixed-grid magnetohydrodynamics (MHD) codes designed to study protoplanetary discs or convection, respectively. Other codes such as Athena++ (Stone et al., 2020) and the new revision K-ATHENA++ (Grete et al., 2020) or DYABLO (Delorme et al., 2022) implement an adaptive mesh refinement (AMR) framework for problems ranging from solar physics to cosmology.
@@
Some codes are more specifically dedicated to the study of core-collapse SNe (for a review, see Müller 2020). These codes were developed initially in 2-D and in the last decade extended to 3-D with different strategies and assets. Their aim is to explore the neutrino-driven explosion mechanism of massive stars with a variety of physical modules (e.g., neutrino transport schemes, neutrino opacities, nuclear networks, general relativity) and spatial grids (e.g. spherical polar, dendritic, Yin-Yang, Cartesian with AMR). These codes include PROMETHEUS and variants (Fryxell et al., 1991; Mueller et al., 1991; Rampp & Janka, 2002; Wongwathanarat et al., 2010, 2013, 2015, 2017; Bollig et al., 2021; Gabler et al., 2021), CHIMERA (Bruenn et al., 2020; Mezzacappa et al., 2020), FORNAX (Skinner et al., 2019; Burrows et al., 2019; Vartanyan et al., 2019; Burrows et al., 2020), COCONOT-FMT (Müller et al., 2019) or FLASH (Couch et al., 2015; O’Connor & Couch, 2018). TESS and variants (Duffell & MacFadyen, 2011, 2013; Duffell, 2016; Mandal & Duffell, 2023) use a variety of expanding grids (moving Voronoi or Cartesian meshes) to study the ejecta as they turn into SN remnants at late times.
@@
In our group, we have developed and used the radiation-hydrodynamics code HERACLES (González et al., 2007) for the study of SN ejecta. The focus has not been the explosion mechanism but instead the evolution of the SN ejecta on a longer timescale. We have studied in 3-D the impact of ejecta structure (i.e., clumping) on Type II SN light curves (Dessart & Audit, 2019). We have also studied the interaction of SN ejecta with circumstellar material, both in spherical symmetry (Dessart et al., 2015) and in 2-D (Vlasis et al., 2016). One asset of our approach is the post-processing of the radiation-hydrodynamics simulations with detailed nonlocal thermodynamic equilibrium (NLTE) radiative transfer for the production of not just light curves but also spectra with CMFGEN (Hillier & Miller, 1998; Dessart & Hillier, 2005; Hillier & Dessart, 2012; Dessart et al., 2015, 2017). Here, we present our current effort at switching from our previous FORTRAN-based CPU-based version of HERACLES to the new code HERACLES++, written in C++, making use of the Kokkos library, and adapted for exascale architectures.
@@
In this paper, we first present the hydrodynamic equations solved for in HERACLES++ as well as the physics treated (§2). Then we describe their implementation and the numerical scheme used for the resolution of the equations (§3). Some benchmarking tests are presented in §4, followed by a more astrophysics-oriented test in §5 with the modeling of a RSG explosion from 500 s after the onset of the explosion until shock breakout. We present our conclusions in §6. We also include several appendix sections to present the results from resolution tests (§A) , performance tests (§B) , as well as additional benchmarking tests (§C).
@@ 2 Formulation of hydrodynamical equations and physics
In this initial version of HERACLES++, we solve the equations of hydrodynamics and ignore radiative transfer. We include various equations of state (EoS), gravity, as well as the possibility for a local heating/cooling term. The code is designed to work in Cartesian or spherical coordinates, and in 1-D, 2-D or 3-D.
@@ 2.1 Hydrodynamics
The equations of hydrodynamics in their conservative form with gravity are given by:
@@
$\frac{\partial(\rho f_{\textsf{x}})}{\partial t}+\vec{\nabla}\cdot(\rho\vec{u}f_{\textsf{x}})$ $=0$ (1d)
where $\rho$ is the density, $\vec{u}$ the velocity, $\mathbb{P}$ the pressure tensor, $P$ the pressure and ${g}$ the gravitational acceleration. The total energy per unit volume $E$ is the sum of the kinetic energy $e_{\rm k}=\rho u^{2}/2$ and the internal energy $e$. The user-defined, power source term $\dot{e}_{\rm hc}$ is used to handle a heating or a cooling process.
The last equation (Eq. 1d) can be added to track the distribution of elements in the course of the simulation. $f_{\textsf{x}}$ is the mass fraction of species x and defined such that $\sum_{\textsf{x}}f_{\textsf{x}}=1$.
@@ 2.2 Gravity terms
Different types of gravity are implemented in HERACLES++. The first option is uniform gravity, which is defined by:
$\vec{g}=\vec{g_{0}}$ (2)
@@
A second option treats the gravity from a central mass $M_{\rm c}$ in spherical geometry. The gravitational acceleration is then given by:
$\vec{g}=\vec{g}(r)=-\frac{GM_{\rm c}}{r^{2}}\,\vec{e_{r}}$ (3)
@@
Self gravity of the outer layers can also be accounted for in the approximation of a spherically symmetric matter distribution. In this case we have:
$\vec{g}=\vec{g}(r,t)=-\frac{GM(r,t)}{r^{2}}\,\vec{e_{r}}$ (4)
@@ 2.3 Equation of state
HERACLES++ can work with a generic EoS. In the present work, two EoS are used. The first one treats the ideal gas where the pressure of the gas and the sound speed are given by:
$P=P_{g}=(\gamma-1)e\quad\textrm{ and}\quad c_{s}=\sqrt{\frac{\gamma\,P_{g}}{\rho}}$ (5)
@@
The second EoS treats together an ideal gas in equilibrium with radiation. In this case the total internal energy and pressure are the sum of the gas and the radiative components:
@@
$P=P_{g}+P_{r}\quad\textrm{with}\quad P_{r}=\frac{1}{3}\,a_{r}T^{4}$ (6b)
where $k_{b}$ is the Boltzman constant, $\mu$ the mean molecular weight and $a_{r}$ is the radiative constant. The sound speed is then given by :
$c_{s}=\sqrt{\frac{\gamma/(\gamma-1)+20\alpha+16\alpha^{2}}{1/(\gamma-1)+12\alpha}\frac{P_{g}}{\rho}}$ (7)
@@
For this EoS, the pressure cannot be computed directly from the internal energy, so we first need to determine the temperature. This is done using the internal energy equation 6a and a Newton-Raphson approach, which converges in approximately ten iterations. This EoS is used for simulations of stellar explosions, which are characterized by radiation-dominated conditions. This also allows for the continuation with radiation hydrodynamics of an originally hydrodynamics-only simulation (the switch should be done when somewhere on the grid the conditions are ripe for photon decoupling from the gas).
@@ 2.4 Energy source terms: heating and cooling
A term $\dot{e}_{\rm hc}$ may be added to the energy equation (Eq. 1c) in order to study specific astrophysical conditions. It may be positive (heating) or negative (cooling). It allows for example to treat radioactive decay or radiative losses. The user can define a custom function fitted to the physical setup being studied.
When studying SN ejecta, unstable isotopes decay and release neutrinos, $\gamma$ rays, and positrons in the process. At present, we treat the ${}^{56}\rm{N}$i decay chain, with daughter isotopes ${}^{56}\rm{C}$o and ${}^{56}\rm{F}$(Nadyozhin, 1994). The associated term in the energy equation is given by:
$\dot{e}_{\rm hc}(t)=\lambda\Big{\{}\Big{[}Q_{\rm Ni}\Big{(}\frac{\tau_{\rm Co}}{\tau_{\rm Ni}}-1\Big{)}-Q_{\rm Co}\Big{]}e^{-t/\tau_{\rm Ni}}+Q_{\rm Co}e^{-t/\tau_{\rm Co}}\Big{\}}\,\rho f_{{\rm Ni}}$ (8)
where $Q_{\rm Ni}=1.75$ MeV and $Q_{\rm Co}=3.73$ MeV are the total energy ($\gamma$-ray and positrons) emitted per decay of ${}^{56}\rm{N}$i and ${}^{56}\rm{C}$o, and $\lambda$ is given by :
@@
The lifetime of ${}^{56}\rm{N}$i and ${}^{56}\rm{C}$o are given by $\tau_{\rm Ni}=$ 8.8 d and $\tau_{\rm Co}=$ 111.3 d. At the present time, we assume that this decay power is fully absorbed locally.
For radioactive decay heating, this emissivity term depends only the total yield of ${}^{56}\rm{N}$i and its spatial distribution (i.e., the total number of ${}^{56}\rm{N}$i nuclei at all radii). However, $\dot{e}_{\rm hc}$ may also depend on temperature, as in the cases of nuclear burning or radiative cooling and in this case one must iterate to find the proper $\dot{e}_{\rm hc}$ and temperature for the corresponding EoS.

### v2

# HERACLES++: A multidimensional Eulerian code for exascale computing

@@
HERACLES++: A multidimensional Eulerian code for exascale computing
@@
Numerical simulations of multidimensional astrophysical fluids present considerable challenges. However, the development of exascale computing has significantly enhanced computational capabilities, motivating the development of new codes that can take full advantage of these resources. In this article, we introduce HERACLES++, a new hydrodynamics code with high portability, optimized for exascale machines with different architectures and running efficiently both on CPUs and GPUs. The code is Eulerian and employs a Godunov finite-volume method to solve the hydrodynamics equations, which ensures accuracy in capturing shocks and discontinuities. It includes different Riemann solvers, equations of state, and gravity solvers. It works in Cartesian and spherical coordinates, either in 1D, 2D, or 3D, and uses passive scalars to handle gases with several species. The code accepts a user-supplied heating or cooling term to treat a variety of astrophysical contexts. In addition to the usual series of benchmarking tests, we use HERACLES++ to simulate the propagation of a supernova shock in a red-supergiant star envelope, from minutes after core collapse until shock emergence. In 1D, the results from HERACLES++ are in agreement with those of V1D for the same configuration. In 3D, the Rayleigh-Taylor instability develops and modifies the 1D picture by introducing density and composition fluctuations as well as turbulence. The focus on a wedge, rather than the full solid angle, and the ability to run HERACLES++ with a large number of GPUs allow for long-term simulations of 3D supernova ejecta with a sub-degree resolution. Future developments will extend HERACLES++ to a radiation-hydrodynamics code.
@@ 1 Introduction
Massive stars end their lives in the gravitational collapse of their degenerate core and the formation of a compact object, possibly followed by an energetic explosion. This phenomenon is dubbed a core-collapse supernova (SN; Baade & Zwicky 1934). Most of the binding energy released by the collapsing envelope is emitted as neutrinos (Arnett et al., 1989). If revived, the SN shock born at core bounce propagates through the progenitor envelope and its highly stratified composition structure. During its expansion through the progenitor envelope, hydrodynamical instabilities grow, leading to chemical mixing on small and large scales. Spatially resolved SN remnants provide direct evidence for this chemical mixing, as well as the heterogeneous density structure of SN ejecta, with well-documented examples such as SN 1987A or Cas A (Fesen et al., 2006; Abellán et al., 2017; Milisavljevic et al., 2024). In type II SNe, which arise from the explosion of generally red-supergiant (RSG) stars, the main instability is Rayleigh-Taylor (Falk & Arnett, 1973). Its development is influenced by the post-shock neutrino-driven convection that develops as part of the explosion (Kifonidis et al., 2000, 2003), as well as the structured convective envelopes of these extended progenitor stars (Goldberg et al., 2022a, b). The nonlinear growth of the Rayleigh-Taylor instability in SN ejecta has been extensively studied (e.g., see Chevalier 1976) or more recently Duffell 2016. Most of these studies were initially focused on SN 1987A (Ebisuzaki et al., 1989; Mueller et al., 1991; Fryxell et al., 1991; Hachisu et al., 1994) and have been extended in recent years to type II SNe in general (e.g., Wongwathanarat et al. 2015).
@@
When studying the development of instabilities in SN ejecta, the structure and chemical composition of the progenitor is critical. The Rayleigh-Taylor instability tends to develop at the H-He and the O-He interfaces (Herant & Benz, 1991; Mueller et al., 1991; Herant & Woosley, 1993), and the resulting structures take the form of fingers and mushrooms, altering the density and chemical profiles obtained in spherical symmetry (e.g., Wongwathanarat et al. 2015). The modeling of such fluids in 3D is computationally intensive and it is therefore desirable to develop efficient and robust numerical tools to tackle these scientific problems.
@@
More and more computational capacities are available with the creation of new supercomputers and the development of new hardware. Exascale supercomputers operate with a massively parallel architecture and have a computing power exceeding $10^{18}$ floating-point operations per second (FLOPS). The exascale era offers the ability to solve large and complex problems, opening up new possibilities for research by combining unprecedented computing power and data analysis. Indeed, the computational resources rely less and less on CPUs and benefit more and more from GPUs. Practically, among the world’s fastest supercomputers, more than 75% of the available FLOPS are delivered by GPUs [^1] Computed from www.top500.org. GPUs have become a crucial driver of computational performance in high-performance computing (HPC) and different vendors (e.g., AMD, NVIDIA, Intel) are developing their own types of hardware, which can be used with their own programming languages (e.g., HIP/ROCm, CUDA, OneApi). To fully exploit these possibilities, scientists must develop software and numerical approaches that use the benefits of exascale hardware and ensure portability across different architectures.
@@
To meet these challenges, and especially performance portability, we have developed the hydrodynamics code HERACLES++ using the Kokkos library (Trott et al., 2022). The unified C++ code is portable to different architectures, offering a level of abstraction that enhances user-friendliness. The code uses the Message Passing Interface (MPI) to deal with distributed memory parallelism and Kokkos for shared memory computations. With this approach, the code can efficiently use most of the CPUs and GPUs available on the market, including NVIDIAs and AMDs.
@@
The usage of Kokkos is slowly spreading and different codes have been recently developed using the Kokkos-C++ framework. For example, IDEFIX (Lesur et al., 2023) and ARK (Padioleau et al., 2019; Bourgeois et al., 2024) are fixed-grid magnetohydrodynamics (MHD) codes designed to study protoplanetary discs or convection, respectively. Other codes, such as Athena++ (Stone et al., 2020) and the new revision, K-ATHENA++ (Grete et al., 2020), or DYABLO (Delorme et al., 2022), implement an adaptive mesh refinement (AMR) framework for problems ranging from solar physics to cosmology.
@@
Some codes are more specifically dedicated to the study of core-collapse SNe (for a review, see Müller 2020). These codes were developed initially in 2D and in the last decade extended to 3D with different strategies and assets. Their aim is to explore the neutrino-driven explosion mechanism of massive stars with a variety of physical modules (e.g., neutrino transport schemes, neutrino opacities, nuclear networks, general relativity) and spatial grids (e.g. spherical polar, dendritic, Yin-Yang, Cartesian with AMR). These codes include PROMETHEUS and variants (Fryxell et al., 1991; Mueller et al., 1991; Rampp & Janka, 2002; Wongwathanarat et al., 2010, 2013, 2015, 2017; Bollig et al., 2021; Gabler et al., 2021), CHIMERA (Bruenn et al., 2020; Mezzacappa et al., 2020), FORNAX (Skinner et al., 2019; Burrows et al., 2019; Vartanyan et al., 2019; Burrows et al., 2020), COCONOT-FMT (Müller et al., 2019), or FLASH (Couch et al., 2015; O’Connor & Couch, 2018). TESS and variants (Duffell & MacFadyen, 2011, 2013; Duffell, 2016; Mandal & Duffell, 2023) use a variety of expanding grids (moving Voronoi or Cartesian meshes) to study the ejecta as they turn into SN remnants at late times.
@@
In our group, we have developed and used the radiation-hydrodynamics code HERACLES (González et al., 2007) for the study of SN ejecta. The focus has not been the explosion mechanism but instead the evolution of the SN ejecta on a longer timescale. We have studied in 3D the impact of ejecta structure (i.e., clumping) on type II SN light curves (Dessart & Audit, 2019). We have also studied the interaction of SN ejecta with circumstellar material, both in spherical symmetry (Dessart et al., 2015) and in 2D (Vlasis et al., 2016). One asset of our approach is the post-processing of the radiation-hydrodynamics simulations with detailed nonlocal thermodynamic equilibrium (NLTE) radiative transfer for the production of not just light curves but also spectra with CMFGEN (Hillier & Miller, 1998; Dessart & Hillier, 2005; Hillier & Dessart, 2012; Dessart et al., 2015, 2017). Here, we present our current effort at switching from our previous FORTRAN-based CPU-based version of HERACLES to the new code HERACLES++, written in C++, making use of the Kokkos library, and adapted for exascale architectures.
@@
In this paper, we first present the hydrodynamic equations solved for in HERACLES++ as well as the physics treated (Sect. 2). We then describe their implementation and the numerical scheme used for the resolution of the equations (Sect. 3). Some benchmarking tests are presented in Sect. 4, followed by a more astrophysics-oriented test in Sect. 5 with the modeling of a RSG explosion from 500 s after the onset of the explosion until shock breakout. We present our conclusions in Sect. 6. We also include several appendix sections to present the results from resolution tests (Appendix A), performance tests (Appendix B), as well as additional benchmarking tests (Appendix C).
@@ 2 Formulation of hydrodynamical equations and physics
In this initial version of HERACLES++, we solve the equations of hydrodynamics and ignore radiative transfer. We include various equations of state (EoSs), gravity, as well as the possibility of a local heating and/or cooling term. The code is designed to work in Cartesian or spherical coordinates, and in 1D, 2D, or 3D.
@@ 2.1 Hydrodynamics
The equations of hydrodynamics in their conservative form with gravity are given by
@@
$\frac{\partial(\rho f_{\textsf{x}})}{\partial t}+\vec{\nabla}\cdot(\rho\vec{u}f_{\textsf{x}})$ $=0,$ (1d)
where $\rho$ is the density, $\vec{u}$ the velocity, $\mathbb{P}$ the pressure tensor, $P$ the pressure, and ${g}$ the gravitational acceleration. The total energy per unit volume, $E$, is the sum of the kinetic energy, $e_{\rm k}=\rho u^{2}/2$, and the internal energy, $e$. The user-defined, power source term $\dot{e}_{\rm hc}$ is used to handle a heating or a cooling process. The last equation (Eq. 1d) can be added to track the distribution of elements in the course of the simulation. $f_{\textsf{x}}$ is the mass fraction of species x and defined such that $\sum_{\textsf{x}}f_{\textsf{x}}=1$.
@@ 2.2 Gravity terms
Different types of gravity are implemented in HERACLES++. The first option is uniform gravity, which is defined by
$\vec{g}=\vec{g_{0}},$ (2)
@@
A second option treats the gravity from a central mass, $M_{\rm c}$, in spherical geometry. The gravitational acceleration is then given by
$\vec{g}=\vec{g}(r)=-\frac{GM_{\rm c}}{r^{2}}\,\vec{e_{r}},$ (3)
@@
Self-gravity of the outer layers can also be accounted for in the approximation of a spherically symmetric matter distribution. In this case, we have:
$\vec{g}=\vec{g}(r,t)=-\frac{GM(r,t)}{r^{2}}\,\vec{e_{r}},$ (4)
@@ 2.3 Equation of state
HERACLES++ can work with a generic EoS. In the present work, two EoSs are used. The first one treats the ideal gas, for which the pressure of the gas and the sound speed are given by
$P=P_{g}=(\gamma-1)e\quad\textrm{ and}\quad c_{s}=\sqrt{\frac{\gamma\,P_{g}}{\rho}},$ (5)
@@
The second EoS treats together an ideal gas in equilibrium with radiation. In this case, the total internal energy and pressure are the sum of the gas and the radiative components:
@@
$P=P_{g}+P_{r}\quad\textrm{with}\quad P_{r}=\frac{1}{3}\,a_{r}T^{4},$ (6b)
where $k_{b}$ is the Boltzman constant, $\mu$ the mean molecular weight, and $a_{r}$ is the radiative constant. The sound speed is then given by
$c_{s}=\sqrt{\frac{\gamma/(\gamma-1)+20\alpha+16\alpha^{2}}{1/(\gamma-1)+12\alpha}\frac{P_{g}}{\rho}},$ (7)
@@
For this EoS, the pressure cannot be computed directly from the internal energy, so we first need to determine the temperature. This is done using the internal energy equation 6a and a Newton-Raphson approach, which converges in approximately ten iterations. This EoS is used for simulations of stellar explosions, which are characterized by radiation-dominated conditions. This also enables the continuation with radiation hydrodynamics of an originally hydrodynamics-only simulation (the switch should be done when somewhere on the grid the conditions are ripe for photon decoupling from the gas).
@@ 2.4 Energy source terms: Heating and cooling
A term, $\dot{e}_{\rm hc}$, may be added to the energy equation (Eq. 1c) in order to study specific astrophysical conditions. It may be positive (heating) or negative (cooling). It allows one for example to treat radioactive decay or radiative losses. The user can define a custom function fit to the physical setup being studied.
When studying SN ejecta, unstable isotopes decay and release neutrinos, $\gamma$ rays, and positrons in the process. At present, we treat the ${}^{56}\rm{N}$i decay chain, with daughter isotopes ${}^{56}\rm{C}$o and ${}^{56}\rm{F}$(Nadyozhin, 1994). The associated term in the energy equation is given by
$\dot{e}_{\rm hc}(t)=\lambda\Big{\{}\Big{[}Q_{\rm Ni}\Big{(}\frac{\tau_{\rm Co}}{\tau_{\rm Ni}}-1\Big{)}-Q_{\rm Co}\Big{]}e^{-t/\tau_{\rm Ni}}+Q_{\rm Co}e^{-t/\tau_{\rm Co}}\Big{\}}\,\rho f_{{\rm Ni}},$ (8)
where $Q_{\rm Ni}=1.75$ MeV and $Q_{\rm Co}=3.73$ MeV are the total energy ($\gamma$-ray and positrons) emitted per decay of ${}^{56}\rm{N}$i and ${}^{56}\rm{C}$o, and $\lambda$ is given by
@@
The lifetimes of ${}^{56}\rm{N}$i and ${}^{56}\rm{C}$o are given by $\tau_{\rm Ni}=$ 8.8 d and $\tau_{\rm Co}=$ 111.3 d. At the present time, we assume that this decay power is fully absorbed locally.
For radioactive decay heating, this emissivity term depends only the total yield of ${}^{56}\rm{N}$i and its spatial distribution (i.e., the total number of ${}^{56}\rm{N}$i nuclei at all radii). However, $\dot{e}_{\rm hc}$ may also depend on the temperature, as in the cases of nuclear burning or radiative cooling, and in this case we must iterate to find the proper $\dot{e}_{\rm hc}$ and temperature for the corresponding EoS.